OpenStudy (anonymous):

can someone please help me with this problem?Solve the absoulute value equation I x-2 I=5

OpenStudy (anonymous):

x-2=5 or x-2=-5 so x=7 or -3

OpenStudy (anonymous):

so how do i write the answer

OpenStudy (anonymous):

um x=7,x=3 or x=3,7

OpenStudy (anonymous):

i mean -3

OpenStudy (anonymous):

so i can write x=7 , x=-3

OpenStudy (anonymous):

yeah i think so

OpenStudy (anonymous):

can i ask one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

I x-2 I=3

OpenStudy (anonymous):

ok so x-2=3 or x-2=-3 so x=5,x=-1

OpenStudy (anonymous):

x=5 x=-1 RIGHT lol

OpenStudy (anonymous):

is that right?^

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (anonymous):

no problem